OpenStudy (anonymous):

Which line is parallel to the line y = 3x + 6? (Points : 1) y = -3x – 6 y = 6x + 3 y = -1/3 x + 6 y = 3x – 9 Question 2. 2. Find the equation of a line that is parallel to y = 2x + 3 and passes through (-1, -1). (Points : 1) y = 2x + 1 y= 2x + 3 y = 4x + 3 y = 4x + 5 Question 3. 3. Find the equation of a line that is parallel to y = 2x - 10 and passes through the point (3, -2). (Points : 1) y = 2x – 8 y = 2x + 8 y = -2x – 4 y = -2x

OpenStudy (anonymous):

please help me please i will return the fav

Nnesha (nnesha):

parallel lines have same slope

Nnesha (nnesha):

can you post one at aa timie

OpenStudy (anonymous):

so it would be a

OpenStudy (anonymous):

can you help me with more please

OpenStudy (anonymous):

it has the same slop 3x

OpenStudy (shadowlegendx):

Parallel lines have the same slope by different y intercepts c;

OpenStudy (anonymous):

thats why i said a

Nnesha (nnesha):

A have -3x

OpenStudy (anonymous):

i got that thats a

Nnesha (nnesha):

i'm doing question 1

OpenStudy (shadowlegendx):

Negative 3x Positive 3x Not the same slope

Nnesha (nnesha):

if you post question 1 and the another that will be \(\color{red}{Awesome}\)

OpenStudy (anonymous):

that is what i am doing i said it is Y=-3x+6

Nnesha (nnesha):

then ****

Nnesha (nnesha):

in your question you have y = 3x + 6 what is your slope in this line ???

OpenStudy (anonymous):

Which line is parallel to the line y = 3x + 6? (Points : 1) y = -3x – 6 y = 6x + 3 y = -1/3 x + 6 y = 3x – 9 \

OpenStudy (anonymous):

3x

Nnesha (nnesha):

not X 3 is a slope so parallel lines have same slope so which answer choice have same slope

OpenStudy (anonymous):

b

OpenStudy (anonymous):

d i mean

Nnesha (nnesha):

yes right D

OpenStudy (anonymous):

can you please stay an help me pleae i have 2 ore test that is only 5 questions befor 12 a clock i need help please

Nnesha (nnesha):

\(\color{blue}{Test}\) ??? sorry test question not allowed to ask

OpenStudy (anonymous):

this is the study guide for them

Nnesha (nnesha):

ohh ok

OpenStudy (anonymous):

if i do not get these turn in tommrw when i take my test i will get a 60 no matter how i good i do on the answer of the test thats why i need help so bad

OpenStudy (anonymous):

so please .......................Question 2. 2. Find the equation of a line that is parallel to y = 2x + 3 and passes through (-1, -1). (Points : 1)

Nnesha (nnesha):

same question parallel lines have same slope

Nnesha (nnesha):

so whats slope of your line gonna be

OpenStudy (anonymous):

would it be a

Nnesha (nnesha):

how can you say that B also have same slope

OpenStudy (anonymous):

becuase be would be on top of each other

Nnesha (nnesha):

yes A is right bec when you solve for b to find y-intercept so plug your order pair into the equation to find y-intercept so yes A is right